OpenStudy (anonymous):

Simplify the rational expressions. State any excluded values. Show your work. x-2/x^2+3x-10

OpenStudy (anonymous):

x^2+3x-10=(x-2)(x+5) now you can simplify but remember you can't Simplify when x=2 and x=-5

OpenStudy (amorfide):

factorise the bottom of the equation \[x^{2}+3x-10=(x+5)(x-2)\] this gives \[\frac{ (x-2) }{ (x+5)(x-2) }\] \[\frac{ x-2 }{ x-2 }=1\] this gives \[\frac{ 1 }{ x+5 }\] \[ x \neq -5\] as it would make you divide by 0

OpenStudy (anonymous):

thanks